two boats leave a port at the same time,one traveling north and the other traveling south.the northbound boat travels at 12 miles per hour faster then the southbound boat.if the southbound boat is traveling at 25 mph,how long before they are 155 miles apart?
Answer by richwmiller(17219)

s=25
n=25+12=37
25t+37t=155
(25+37)*t=155
62t=155
t=2.5 hours